OpenStudy (anonymous):

ok, really starting to get frustrated with math......been trying to finish this section for 2 days now but i keep running into problems.... Find equations of the tangent plane and normal line to the surface x=3y2+4z2−211 at the point (-3, 2, 7). Tangent Plane: (make the coefficient of x equal to 1). Normal line: ⟨−3, , ⟩ +t⟨1, , ⟩.

OpenStudy (anonymous):

i got grad f =<-1, 6y, 8z> and grad f at the point =<-1,12,56> so should be <-1, 12, 56>*<x+3, y-2, z-7> or -1(x+3)+12(y-2)+56(z-7). and of course, webwork says it's wrong.............

OpenStudy (swissgirl):

hi

OpenStudy (swissgirl):

since they want coeff of x = 1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

why are you finding distance? i'm confused.....what's d?

OpenStudy (swissgirl):

huh

OpenStudy (swissgirl):

equation of a plane is given by ax+by+cz=d where a,b,c,d are constants

OpenStudy (swissgirl):

also if u wanna arrive at a relationship u are missing an equality here

OpenStudy (swissgirl):

<-1, 12, 56>*<x+3, y-2, z-7> =0 the dot of these 2 vectors for any given point x,y,z on the plane is 0 since the normal line is perpendicular or "normal" to the plane

OpenStudy (swissgirl):

-1(x+3)+12(y-2)+56(z-7)=0 simply this -x-3+12y-24+56z-56*7=0 so -x+12y+56z=3+24+56*7 x-12y-56z=-(3+24+56*7)

OpenStudy (anonymous):

why can't i use grad f (x0, y0, z0)*(x-x0, y-y0, z-z0>?

OpenStudy (anonymous):

cause that -1(x+3)+12(y-2)+56(z-7)=0 is exactly what i got and it's wrong.

OpenStudy (swissgirl):

looks like i made a silly error the first time it shud be the coeff from the normal so -x+12y+56z=d -3+12(2)+56*7=d x-12y-56z=-(-3+12(2)+56*7)

OpenStudy (anonymous):

i think -1(x+3)+12(y-2)+56(z-7)=0 is right, they just want the coefficient of x to be one, so i don't know how to change that.

OpenStudy (swissgirl):

i just did it for u

OpenStudy (swissgirl):

x-12y-56z=-413

OpenStudy (anonymous):

but that's still not the right answer

OpenStudy (swissgirl):

do they want it entered in some particular form?

OpenStudy (anonymous):

sorry, forgot to make the screen shot .jpeg....gimme a sec

OpenStudy (swissgirl):

oh wait its not -3 hters its --3 so

OpenStudy (swissgirl):

just input it in like that it will accept it its webwork

OpenStudy (swissgirl):

oh =0 so ya x-12y-56z+(-3+12(2)+56*7)=0

OpenStudy (swissgirl):

did it work?

OpenStudy (anonymous):

nope

OpenStudy (swissgirl):

its not 413 u pt anymore right

OpenStudy (anonymous):

Entered Answer Preview Result Messages x-12*y-56*z+(-3)+12*2+56*7 x−12y−56z+(−3)+12⋅2+56⋅7 x-12*y-56*z+(-3)+12*2+56*7 incorrect This answer is equivalent to the one you just submitted.

OpenStudy (anonymous):

hence why i'm really starting to get annoyed with math.....

OpenStudy (swissgirl):

oh change the 3 to -3

OpenStudy (anonymous):

wanna just give up and get an F on this assignment.......

OpenStudy (swissgirl):

wtf.. dude its easy.. change the 3 to -3

OpenStudy (anonymous):

well, like i said in the beginning of the post, i keep running into problems. this is only qeustion 11 of 20 and i've run into issues on most all of them

OpenStudy (swissgirl):

x-12y-56z+(-(-3)+12(2)+56*7)=0

OpenStudy (anonymous):

every time i think i understand something, i'm wrong.

OpenStudy (anonymous):

i didn't know about setting it equal to d....had no idea what you were doing in the beginning.

OpenStudy (swissgirl):

is it right yet

OpenStudy (anonymous):

the first part, finally.

OpenStudy (swissgirl):

...

OpenStudy (swissgirl):

its not my fault.. i dont want to keep scrolling up to see the point... no time to memorize the point was -3 not and 3 for x

OpenStudy (swissgirl):

as far as the equation for d its straight forward

OpenStudy (anonymous):

i didn't say it was your fault. you're helping.

OpenStudy (swissgirl):

you know the dot of normal * your <x-a,y-b...> vectors will leave the normal line's components as the components on x y and z

OpenStudy (anonymous):

just saying i'm extremely frustrated is all. been working on this homework set for 2 days....

OpenStudy (swissgirl):

instead of doing the dot product which is a bit annoying, u can just calculate the normal and u aleady know like if the normal is <a,b,c> then u can say ax+by+cz=d now u plug in the point x,y,z and see what d has to be

OpenStudy (swissgirl):

such that the equation of your plane is satisfied

OpenStudy (swissgirl):

k do the nromal line since they want it at -3 for x

OpenStudy (swissgirl):

and parametrized at 1 change in x

OpenStudy (swissgirl):

we can see what the other vaues have to be

OpenStudy (anonymous):

ax+by+cz=d.....(-3)-12(2)-56(7)+419=d

OpenStudy (swissgirl):

your normal line vector is <-1,12,56> same as <1,-12,-56> right

OpenStudy (swissgirl):

so right away u know this part <3,...,...> +t<1,-12,-56>

OpenStudy (anonymous):

no....that = 0

OpenStudy (swissgirl):

now to see what the starting point must be

OpenStudy (swissgirl):

<-3*,...,...> +t<1,-12,-56>

OpenStudy (anonymous):

ax+by+cz=d.....(-3)-12(2)-56(7)+419=d d= 0 .........which doesn't sound helpful at all........grr......

OpenStudy (swissgirl):

<-3*,12*3,3(56)> +t<1,-12,-56>

OpenStudy (swissgirl):

see if that works

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Entered Answer Preview Result x-12*y-56*z+3+12*2+56*7 x−12y−56z+3+12⋅2+56⋅7 correct 36 12⋅3 incorrect 168 3⋅56 incorrect -12 −12 correct -56 −56 correct

OpenStudy (aum):

In your very first reply you have the equation of the tangent plane correct: -1(x+3)+12(y-2)+56(z-7) = 0 But they want the x-coefficient to be 1. So multiply throughout by -1: (x+3) - 12(y-2) - 56(z-7) = 0 x - 12y - 56z + 419 = 0

OpenStudy (swissgirl):

im asking for the 2nd part

OpenStudy (swissgirl):

can u take screen shot that is confusing to look at

OpenStudy (anonymous):

that was the second part....you went really fast, swissgirl. i'm really lost now. i have no idea what you did.....

OpenStudy (aum):

Normal line = <-3, 2, 7> + t<-1,12,56>

OpenStudy (swissgirl):

oh oops just put the point in <1,

OpenStudy (swissgirl):

for a second i thought they wanted u to just parametrize a line going thru the origin

OpenStudy (swissgirl):

=]

OpenStudy (anonymous):

OH! yeah, cause <x0, y0, z0>+t,...

OpenStudy (anonymous):

my frustration is confusing me and my ADHD is not helping. my brain hurts.

OpenStudy (swissgirl):

alright go on next question

OpenStudy (aum):

Try: <-3, 2, 7> + t<1,-12,-56>

OpenStudy (swissgirl):

he already got it aum

OpenStudy (aum):

Oh, ok.

OpenStudy (swissgirl):

cant u see he doesnt need ur shiet right now the man is stressed!

OpenStudy (swissgirl):

he's stressing me out too

OpenStudy (anonymous):

wait, what? aum helped and you helped....who's stressing who out? math is stressing me. that's all. i got the answer right.

OpenStudy (swissgirl):

just jokin, lets goo next question

OpenStudy (anonymous):

oh, ok.

OpenStudy (anonymous):

next one is very similar. Find equations of the tangent plane and normal line to the surface z−16=xeycosz at the point (-16, 0, 0). Tangent Plane: (make the coefficient of z equal to 1). =0. Normal line: ⟨−16, , ⟩ +t⟨ , , 1⟩.

OpenStudy (swissgirl):

take screnshot

OpenStudy (anonymous):

so first step is finding partial x and partial y.

OpenStudy (anonymous):

kk

OpenStudy (swissgirl):

yah

OpenStudy (swissgirl):

u know why the partial is working right

OpenStudy (anonymous):

tangent is derivative or slope. yeah.

OpenStudy (swissgirl):

ok ya do the same thing write it as G=f(x,y,z)

OpenStudy (swissgirl):

<G_x,G_y,G_z>

OpenStudy (anonymous):

i have to get z-16 on the other side, right?

OpenStudy (swissgirl):

not hte 16 we dont care about constants when it comes to the slopes

OpenStudy (swissgirl):

i mean constants that in a term by themself that is

OpenStudy (anonymous):

yeah, but we have to find the partial of z so it has to be on the same side?

OpenStudy (swissgirl):

http://prntscr.com/4u50nl

OpenStudy (anonymous):

yeah, that's wht i got for partials. just a little slower to get there.